OpenStudy (anonymous):

How to find the SUM of the solution of the equation: 8x^2+40x=16x

OpenStudy (anonymous):

8x^2 + 40x = 16x (-16x) 8x^2 + 32x = 0 8x(x + 4) = 0 I'll let you take it from here.

OpenStudy (anonymous):

Set 8x = 0 and (x +4) = 0 , then the results add together

OpenStudy (anonymous):

Correct, since it asks for the sum. You get two results, one from either factor.

OpenStudy (anonymous):

Thanks! How did you get 32x?

OpenStudy (anonymous):

By subtracting 16x from both sides and then grouping like terms. Like this: 8x^2 +40x = 16x 8x^2 +40x -16x = 16x-16x 8x^2 +32x = 0

OpenStudy (anonymous):

So you subtracted 16x-16x, do you also subtract 40x-16x? If so, wouldn't that equal 24x?

OpenStudy (anonymous):

Yes. Haha thank you for catching that.

OpenStudy (anonymous):

Embarassing.

OpenStudy (anonymous):

haha. It's okay!

OpenStudy (anonymous):

So after, do you divide each by 8?

OpenStudy (anonymous):

Yeah. I get 8x^2 + 24x = 0 and you can take out a factor of 8x

OpenStudy (anonymous):

So the answer would be 3?

OpenStudy (anonymous):

Kind of almost.

OpenStudy (anonymous):

You get a factor of (x+3) What x makes that factor 0?

OpenStudy (anonymous):

you can factor out as 4x(2x+6) then set both to zero and you should end up with x= 0 & x = -3 and add them

OpenStudy (anonymous):

So the answer is -3?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

That was easy! Thanks you guys!